# Existence of Friedrich-Wintgen Bound States in the Continuum: Cavity with a Thin Waveguide Opening

### v1

@@ 4.1 Decomposition into real and imaginary components
(39) $\alpha_{0}>0,\ \text{and}\ {\bf i}\alpha_{j}<0\ \text{for\ each}\ j\in\mathbb{Z}^{+}.$
@@ 4.2 First reduction
[…]
$=$ $\int_{R_{h}}\partial_{x_{1}}u_{1}(x_{1},x_{2})\phi_{j}(x_{2})\chi_{j}(x_{1})dx_{2}dx_{1}$
(44) $+\int_{R_{h}}\partial_{x_{2}}u_{1}(x_{1},x_{2})\left(-\frac{h}{j\pi}\eta_{j}(x_{2})\chi^{\prime}_{j}(x_{1})\right)dx_{2}dx_{1},$
@@
(45) $\left\|\sqrt{{\bf i}\alpha_{j}}\phi_{j}\chi_{j}\right\|_{L^{2}(R_{h})}^{2}=$ $\frac{1}{3}\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu},$
(46) $\left\|-\frac{h}{j\pi}\sqrt{{\bf i}\alpha_{j}}\eta_{j}\chi_{j}^{\prime}\right\|_{L^{2}(R_{h})}^{2}=$ $\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu}.$
@@
In the following analysis, we always assume $R_{h}\subset{\Omega_{\mathrm{in}}}$. To reduce the system (40)–(41) into a finite-dimensional form, we adopt the following notation:
@@
Under Condition A.1 and the assumption $h\ll 1$, the operator $\mathbf{B}$ is uniformly bounded in $\ell^{2}$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, we can obtain
(50) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\times\ell^{2}}\leq 1,$
@@
Fix $h>0$ sufficiently small that $\mu<\pi^{2}/h^{2}$ and $R_{h}\subset\Omega_{\mathrm{in}}$. We restrict $\delta$ to a small interval that (31)–(32) are satisfied. Let $\{d_{1,m}\}_{m=M}^{\infty}$ and $\{d_{2,m}\}_{m=M}^{\infty}$ denote sequences in $\ell^{2}$. Then we can obtain:
@@
where the first inequality follows from Lemma 4.3, and the second inequality follows from our definition of $H^{1}(\Omega_{\mathrm{in}})$ norm in (15). Since $-\mathbf{B}$ is positive, it follows that
(53) $((\mathbf{I}-\mathbf{B})\mathbf{v},\mathbf{v})_{\ell^{2}}\geq\|\mathbf{v}\|_{\ell^{2}}^{2},\ \text{for}\ \mathbf{v}\in\ell^{2}.$
@@
(54) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\times\ell^{2}}\|\widetilde{\mathbf{v}}\|^{2}_{\ell^{2}}\geq(\widetilde{\mathbf{v}},(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}})_{\ell^{2}}\geq\|(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}}\|_{\ell^{2}}^{2},$
@@
Consider the subsystem in (40) satisfied by $[d_{M},d_{M+1},\ldots]^{T}$:
@@
Theorem 4.5 implies $[d_{M},d_{M+1},\ldots]^{T}$ can be represented by $[d_{0},\cdots,d_{M-1}]^{T}$, which leads to a finite dimensions linear system:
@@ 4.3 Further reduction
We further reduce the system (56)–(57) to a $3\times 2$ linear system for the vector $[d_{M-2},d_{M-1}]^{T}$. By evaluating the determinants of two submatrices, we derive two governing equations. The existence of BICs corresponds to determining pairs $(\delta,\mu)$ for which both equations are satisfied. To formalize this, we introduce the following notation:
@@
Under Condition A.1 and the assumption $h\ll 1$, each entry in the matrices $\mathbf{C}_{0,0},\ \mathbf{C}_{0,1},\ \mathbf{C}_{1,0}$ and $\mathbf{C}_{1,1}$ converges to 00 uniformly as $h\to 0$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38).
@@
We demonstrate this result for each entry in $\mathbf{C}_{0,0}$. Assume $h\ll 1$ such that $\mu<\pi^{2}/h^{2}$ and $R_{h}\subset\Omega_{\mathrm{in}}$. We restrict $\delta$ to a small interval to ensure the validity of (31)–(32). Let $c_{m_{1},m_{2}}$ denote the entry in the $m_{1}$-th row and $m_{2}$-th column of $\mathbf{C}_{0,0}$. Then $c_{m_{1},m_{2}}$ can be expressed as:
@@
For fixed $h$, consider the estimate:
$\left|\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{1})\right\|_{[L^{2}(R_{h})]^{2}}-\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{2})\right\|_{[L^{2}(R_{h})]^{2}}\right|$
$\leq$ $\left\|\frac{1}{n}\left(\nabla\psi_{m}(\cdot,\delta_{1})-\nabla\psi_{m}(\cdot,\delta_{2})\right)\right\|_{[L^{2}(R_{h})]^{2}}$
(66) $\leq$ $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{\mathrm{in}})}.$
@@
Under Condition A.1 and the assumption $h\ll 1$, the matrix $\mathbf{L}_{0,0}-\mathbf{C}_{0,0}$ is invertible for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, there exists a constant $C_{1}>1/\epsilon$ such that
@@
Restrict $\delta$ to a sufficiently small interval ensuring the validity of (31)–(32). For $\mu$ satisfying (38), the matrix $\mathbf{L}_{0,0}$ is negative definite with eigenvalues less than $-\epsilon$, which guarantees its invertibility. By Lemma 4.7, the operator norm $\|\mathbf{C}_{0,0}\|_{2}\to 0$ as $h\to 0$. Choose $h\ll 1$ such that $\|\mathbf{C}_{0,0}\|_{2}\leq\epsilon-1/C_{1}$. Applying the Neumann series expansion, we obtain:
@@
Utilizing Theorem 4.9, the coefficients $[d_{0},\ldots,d_{M-3}]^{T}$ can be determined from
@@
Here $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ denote the entries of $-\mathbf{C}_{1,1}-\mathbf{C}_{1,0}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}$, while $a_{2,0}$ and $a_{2,1}$ correspond to the coefficients in (70). Thus, the existence of BICs is equivalent to finding solutions to the following equations:
@@ 5 Existence of BICs under analytic medium perturbations
5.1 Uniform convergence of $f_{0}$ and $f_{1}$ to 00 as $h\to 0$
We show $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are well-defined and approach 00 as $h\to 0$ uniformly over small intervals of $\delta$ and $\mu$.
@@ 5.1 Uniform convergence of $f_{0}$ and $f_{1}$ to 00 as $h\to 0$
Let $\Omega_{b}\subset{\Omega_{\mathrm{in}}}$ be a region with a $C^{1}$ boundary satisfying $\overline{\Omega_{b}}\subset\Omega_{B}\cup\{x_{1}=0\}$ and $\Gamma_{h}\subset\partial{\Omega_{b}}$. Under Condition A.1, the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\Omega_{b}}$ is continuous from $[-\delta_{0},\delta_{0}]$ into $H^{3}(\Omega_{b})$ for each $m\in\mathbb{N}$, i.e.,
(76) $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{3}(\Omega_{b})}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}].$
@@
(77) $-\Delta\psi_{m}(\mathbf{x},\delta)=\lambda_{m}(\delta)\psi_{m}(\mathbf{x},\delta),\ \text{for}\ \mathbf{x}\in\Omega_{B}.$
The mapping $\delta\to-\Delta\psi_{m}(\cdot,\delta)|_{\Omega_{B}}$ is therefore continuous from $[-\delta_{0},\delta_{0}]$ into $H^{1}(\Omega_{B})$ by using Lemma 3.1. Since $\partial_{\bm{\nu}}\psi_{m}(\cdot,\delta)=0$ on $\partial\Omega_{\mathrm{in}}$, the regularity theory ensures that $\psi_{m}(\cdot,\delta)\in{H^{3}(\Omega_{b})}$ (cf. Theorem 4.18 in [18]). Furthermore, for $\delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}]$, the estimate
$\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{3}(\Omega_{b})}{\leq}$ $C\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{B})}$
(78) $+C\|\Delta\psi_{m}(\cdot,\delta_{1})-\Delta\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{B})}.$
@@
Under Conditions A.1, A.2, and the assumption $h\ll 1$, the functions $f_{0}$ and $f_{1}$ are well-defined for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, the uniform convergence
(79) $\lim_{h\to 0}f_{0}(\cdot,\cdot,h)=0,\ \text{and}\ \lim_{h\to 0}f_{1}(\cdot,\cdot,h)=0$
hold.
@@
From Lemma 4.7 and Theorem 4.9, the terms $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ are all well-defined and tend to 00 uniformly as $h\to 0$. We focus on showing the uniform boundedness of ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$, and (79) follows. From (70), $a_{2,0}$ can be written as
@@
(80) $+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}(\psi_{M-2},\phi_{j})_{\Gamma_{h}}.$
Here, $\mathbf{q}$ denotes the finite-dimensional vector:
@@
And the following estimate holds by using Lemma 4.3 and Theorem 4.5:
@@
Accordingly, we have
@@
(83) $+n_{1}^{2}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\left\|\frac{1}{n}\nabla\psi_{M-2}\right\|_{[L^{2}(R_{h})]^{2}}\|\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}\|_{\ell^{2}},$
where Theorem 4.9 has been used. Consider the sequence $\{d_{0,m}\}_{m=M}^{\infty}\in\ell^{2}$ and analyze its inner product involving $\mathbf{v}^{*}_{M,\infty,0}\mathbf{D}$. This leads to:
@@
where $C$ is a constant independent of $h$. Consequently, both the norms of $\mathbf{v}_{0,M-3,0}^{*}$ and $\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}$ are uniformly bounded. For $h\ll 1$, the region $R_{h}$ lies within $\Omega_{b}$ as defined in Lemma 5.1. By the general Sobolev inequalities (cf. Section 5 in [3]), (76) gives the continuity of the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\overline{\Omega_{b}}}$ in $C^{1,\gamma}$ for $10<\gamma<$ 1. Thus the following asymptotic expansions hold:
@@
where the constants implicit in the $\mathcal{O}-$notation are independent of $\delta$. From these results, the leading behavior of $a_{2,0}$ is given by:
(89) $a_{2,0}=\psi_{M-2}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h),$
and analogously,
(90) $a_{2,1}=\psi_{M-1}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h),$
which ensures the uniform boundedness of the ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$ for all $\delta$ in a sufficiently small interval.

### v2

@@ 4.1 Decomposition into real and imaginary components
(40) $\alpha_{0}>0\ \text{and}\ {\bf i}\alpha_{j}<0\ \text{for\ every}\ j\in\mathbb{Z}^{+}.$
@@ 4.2 First reduction
$(u_{1},\phi_{j})_{\Gamma_{h}}=$ $\int_{-{h}/{2}}^{{h}/{2}}u_{1}(0,x_{2})\phi_{j}(x_{2})dx_{2}$
$=$ $\frac{h}{j\pi}\int_{R_{h}}\nabla\cdot({u}_{1}(x_{1},x_{2})\mathrm{curl}_{2}\left(\eta_{j}(x_{2})\chi_{j}(x_{1})\right))dx_{1}dx_{2}$
$=$ $\int_{R_{h}}\partial_{x_{1}}u_{1}(x_{1},x_{2})\phi_{j}(x_{2})\chi_{j}(x_{1})dx_{1}dx_{2}$
(45) $+\int_{R_{h}}\partial_{x_{2}}u_{1}(x_{1},x_{2})\left(-\frac{h}{j\pi}\eta_{j}(x_{2})\chi^{\prime}_{j}(x_{1})\right)dx_{1}dx_{2},$
@@
(46) $\left\|\sqrt{{\bf i}\alpha_{j}}\phi_{j}\chi_{j}\right\|_{L^{2}(R_{h})}^{2}=$ $\frac{1}{3}\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu}<1,$
(47) $\left\|-\frac{h}{j\pi}\sqrt{{\bf i}\alpha_{j}}\eta_{j}\chi_{j}^{\prime}\right\|_{L^{2}(R_{h})}^{2}=$ $\sqrt{1-\frac{h^{2}}{\pi^{2}j^{2}}\mu}<1.$
@@
Let $h_{0}>0$ be chosen so that
(48) $\mu_{r}<\lambda_{M}(0)-\epsilon<\frac{\pi^{2}}{h_{0}^{2}}\ \text{and}\ R_{h_{0}}\subset\Omega_{B}.$
In the following analysis, we always assume $h\in(0,h_{0})$. To reduce the system (41)–(42) into a finite-dimensional form, we adopt the following notation:
@@
Under Condition A.1 and the assumption $h\in(0,h_{0})$, the operator $\mathbf{B}$ is uniformly bounded in $\ell^{2}$ for $\delta\in[-\delta_{2},\delta_{2}]$ and $\mu\in[\mu_{l},\mu_{r}]$. Furthermore, we can obtain
(52) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\to\ell^{2}}\leq 1,$
@@
Let $\{d_{1,m}\}_{m=M}^{\infty}$ and $\{d_{2,m}\}_{m=M}^{\infty}$ denote sequences in $\ell^{2}$. A direct computation gives
@@
where the first inequality follows from Lemma 4.3 and the second inequality follows from our definition of $H^{1}(\Omega_{\mathrm{in}})$ norm in (16). Since $\mathbf{B}$ is positive, it follows that
[…]
@@
(56) $\|(\mathbf{I}-\mathbf{B})^{-1}\|_{\ell^{2}\to\ell^{2}}\|\widetilde{\mathbf{v}}\|^{2}_{\ell^{2}}\geq(\widetilde{\mathbf{v}},(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}})_{\ell^{2}}\geq\|(\mathbf{I}-\mathbf{B})^{-1}\widetilde{\mathbf{v}}\|_{\ell^{2}}^{2},$
@@
Consider the subsystem in (41) satisfied by $d_{M},d_{M+1},\ldots]^{T}$:
@@
Theorem 4.5 implies $d_{M},d_{M+1},\ldots]^{T}$ can be represented by $d_{0},\cdots,d_{M-1}]^{T}$, which leads to a finite dimensions linear system:
@@ 4.3 Further reduction
We further reduce the system (58)–(59) to a $3\times 2$ linear system for the vector $d_{M-2},d_{M-1}]^{T}$. By evaluating the determinants of two submatrices, we derive two governing equations. The existence of BICs corresponds to determining pairs $(\delta,\mu)$ for which both equations are satisfied. To formalize this, we introduce the following notation:
@@
Under Condition A.1 and the assumption $h\in(0,h_{0})$, each entry in the matrices $\mathbf{C}_{0,0},\ \mathbf{C}_{0,1},\ \mathbf{C}_{1,0}$ and $\mathbf{C}_{1,1}$ converges to 0 uniformly as $h\to 0$ for $\delta\in[-\delta_{2},\delta_{2}]$ and $\mu\in[\mu_{l},\mu_{r}]$.
@@
We demonstrate this result for each entry in $\mathbf{C}_{0,0}$. Let $c_{m_{1},m_{2}}$ denote the entry in the $m_{1}$-th row and $m_{2}$-th column of $\mathbf{C}_{0,0}$. Then $c_{m_{1},m_{2}}$ can be expressed as:
@@
For fixed $h$, consider the estimate for $\delta_{a},\delta_{b}\in[-\delta_{2},\delta_{2}]$:
$\left|\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{a})\right\|_{[L^{2}(R_{h})]^{2}}-\left\|\frac{1}{n}\nabla\psi_{m}(\cdot,\delta_{b})\right\|_{[L^{2}(R_{h})]^{2}}\right|$
$\leq$ $\left\|\frac{1}{n}\left(\nabla\psi_{m}(\cdot,\delta_{a})-\nabla\psi_{m}(\cdot,\delta_{b})\right)\right\|_{[L^{2}(R_{h})]^{2}}$
(68) $\leq$ $\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{H^{1}(\Omega_{\mathrm{in}})}.$
@@
Under Condition A.1, there exists $h_{1}>0$ such that the matrix $\mathbf{L}_{0,0}-\mathbf{C}_{0,0}$ is invertible for all $h\in(0,h_{1})$, $\delta\in[-\delta_{2},\delta_{2}]$ and $\mu\in[\mu_{l},\mu_{r}]$. Moreover, there exists a constant $C_{1}>1/\epsilon$ such that
@@
Under the assumed ranges of $\delta$ and $\mu$, the matrix $\mathbf{L}_{0,0}$ is negative definite with eigenvalues less than $\epsilon$, ensuring its invertibility. By Lemma 4.7, the operator norm $\|\mathbf{C}_{0,0}\|_{2}\to 0$ as $h\to 0$. Choose $h_{1}<h_{0}$ such that $\|\mathbf{C}_{0,0}\|_{2}\leq\epsilon-1/C_{1}$ for all $h<h_{1}$. Then, by the Neumann series expansion, we obtain:
@@
Utilizing Theorem 4.9, the coefficients $d_{0},\ldots,d_{M-3}]^{T}$ can be determined from
@@
Here $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ denote the entries of $\mathbf{C}_{1,1}-\mathbf{C}_{1,0}(\mathbf{L}_{0,0}-\mathbf{C}_{0,0})^{-1}\mathbf{C}_{0,1}$, while $a_{2,0}$ and $a_{2,1}$ correspond to the coefficients in (72). Thus, the existence of BICs is equivalent to finding solutions to the following equations:
@@ 5 Existence of BICs under analytic medium perturbations
5.1 Uniform convergence of $f_{0}$ and $f_{1}$ to 0 as $h\to 0$
We show $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are well-defined and approach 0 as $h\to 0$ uniformly over small intervals of $\delta$ and $\mu$.
@@ 5.1 Uniform convergence of $f_{0}$ and $f_{1}$ to 0 as $h\to 0$
Let $\Omega_{b}\subset{\Omega_{\mathrm{in}}}$ be a region with a $C^{1}$ boundary satisfying $\overline{\Omega_{b}}\subset\Omega_{B}\cup\{x_{1}=0\}$ and $\Gamma_{h}\subset\partial{\Omega_{b}}$. Under Condition A.1, the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\Omega_{b}}$ is continuous from $\delta_{0},\delta_{0}]$ to $H^{3}(\Omega_{b})$ for each $m\in\mathbb{N}$, i.e.,
(78) $\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{H^{3}(\Omega_{b})}\to 0\ \text{as}\ \delta_{a}\to\delta_{b},\ \text{for}\ \delta_{a},\delta_{b}\in[-\delta_{0},\delta_{0}].$
@@
(79) $\Delta\psi_{m}(\mathbf{x},\delta)=\lambda_{m}(\delta)\psi_{m}(\mathbf{x},\delta),\ \text{for}\ \mathbf{x}\in\Omega_{B}.$
The mapping $\delta\to-\Delta\psi_{m}(\cdot,\delta)|_{\Omega_{B}}$ is therefore continuous from $\delta_{0},\delta_{0}]$ to $H^{1}(\Omega_{B})$ by using Lemma 3.1. Since $\partial_{\bm{\nu}}\psi_{m}(\cdot,\delta)=0$ on $\partial\Omega_{\mathrm{in}}$, the regularity theory ensures that $\psi_{m}(\cdot,\delta)\in{H^{3}(\Omega_{b})}$ (cf. Theorem 4.18 in [20]). Furthermore, for $\delta_{a},\delta_{b}\in[-\delta_{0},\delta_{0}]$, the estimate
$\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{H^{3}(\Omega_{b})}{\leq}$ $C\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{H^{1}(\Omega_{B})}$
(80) $+C\|\Delta\psi_{m}(\cdot,\delta_{a})-\Delta\psi_{m}(\cdot,\delta_{b})\|_{H^{1}(\Omega_{B})}.$
@@
Under Conditions A.1 and A.2, there exist constants $h_{2},\delta_{3}>0$ such that for all $h\in(0,h_{2})$, $\delta\in[-\delta_{3},\delta_{3}]$ and $\mu\in[\mu_{l},\mu_{r}]$, the functions $f_{0}$ and $f_{1}$ are well-defined. Moreover, these functions vanish uniformly in the limit:
(81) $\lim_{h\to 0}f_{0}(\cdot,\cdot,h)=0,\ \text{and}\ \lim_{h\to 0}f_{1}(\cdot,\cdot,h)=0.$
@@
By Lemma 4.7 and Theorem 4.9, the terms $a_{0,0}$, $a_{0,1}$, $a_{1,0}$ and $a_{1,1}$ are all well-defined and converge uniformly to 0 as $h\to 0$. The proof therefore reduces to showing that the ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$ are uniformly bounded, from which (81) follows. From (72), we express $a_{2,0}$ as:
@@
(82) $+\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{D}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}\mathbf{v}_{M,\infty,j}(\psi_{M-2},\phi_{j})_{\Gamma_{h}},$
with the vector $\mathbf{q}$ given by:
@@
An application of Lemma 4.3 and Theorem 4.5 yields the estimate:
@@
Using this and Theorem 4.9, we obtain
@@
(85) $+n_{1}^{2}\sqrt{1+\frac{{\lambda_{M}(0)+1}}{\epsilon}}\left\|\frac{1}{n}\nabla\psi_{M-2}\right\|_{[L^{2}(R_{h})]^{2}}\|\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}\|_{\ell^{2}}.$
Consider the sequence $\{d_{0,m}\}_{m=M}^{\infty}\in\ell^{2}$. The inner product with $\mathbf{v}^{*}_{M,\infty,0}\mathbf{D}$ satisfies:
@@
where $C$ is a constant independent of $h$. Hence, $\mathbf{v}_{0,M-3,0}^{*}$ and $\mathbf{v}_{M,\infty,0}^{*}\mathbf{D}$ are uniformly bounded. For small $h$, $R_{h}$ lies within $\Omega_{b}$ as defined in Lemma 5.1. By the general Sobolev inequalities (cf. Section 5 in [4]), the convergence in (78) implies that the mapping $\delta\to\psi_{m}(\cdot,\delta)|_{\overline{\Omega_{b}}}$ is continuous in the $C^{1,\gamma}$ norm for $0<\gamma<1$. This continuity yields the following asymptotic expansions:
@@
where the constants in the $\mathcal{O}$-terms are independent of $\delta$. From these results, the leading behavior of $a_{2,0}$ is:
(91) $a_{2,0}=\psi_{M-2}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h).$
An analogous argument gives:
(92) $a_{2,1}=\psi_{M-1}(\mathbf{o},\delta)\sqrt{h}+\mathcal{O}(h).$
Condition A.2 implies that
(93) $\frac{\psi_{M-2}(\mathbf{o},0)}{\psi_{M-1}(\mathbf{o},0)}\neq 0\ \text{and}\ \frac{\psi_{M-1}(\mathbf{o},0)}{\psi_{M-2}(\mathbf{o},0)}\neq 0.$
Therefore, we can choose $h_{2},\delta_{3}>0$ such that the ratios ${a_{2,0}}/{a_{2,1}}$ and ${a_{2,1}}/{a_{2,0}}$ are uniformly bounded for all $h\in(0,h_{2})$ and $\delta\in[-\delta_{3},\delta_{3}]$.